𝑓 (𝑛) (𝑧) = 𝑛! /2𝜋𝑖 ∫𝜕𝐷 𝑓(𝑤)𝑑𝑤 / (𝑤−𝑧)𝑛+1 . using the generalized Cauchy integral formula (GCIF) expressed in

Prove it with the Mathematical Induction Method.(As is known, the formula for = 1 is Cauchy Integral

It becomes the Cauchy Integral Formula (CIF), which is the result of the theorem and it is correct.)


1
Expert's answer
2020-08-24T18:38:25-0400

Proof by induction:

For all "n\in\N_{>0}," let "P(n)" be the proposition:


"f^{(n)}(z)=\dfrac{n!}{2\pi i}\int_{\partial D}\dfrac{f(w)}{(w-z)^{n+1}}dw"

Basis for the Induction

"P(0)" holds, as this is:


"f^{(0)}(z)=\dfrac{0!}{2\pi i}\int_{\partial D}\dfrac{f(w)}{(w-z)^{0+1}}dw"

"f(z)=\dfrac{1}{2\pi i}\int_{\partial D}\dfrac{f(w)}{w-z}dw"

which is Cauchy's Integral Formula.

This is our basis for the induxtion. 


Induction Hypothesis

Now we need to show that, if "P(k)" is true, where "k\geq0," then it logically follows that "P(k+1)" is true.

So this is our induction hypothesis:


"f^{(k)}(z)=\dfrac{k!}{2\pi i}\int_{\partial D}\dfrac{f(w)}{(w-z)^{k+1}}dw"

Then we need to show:


"f^{(k+1)}(z)=\dfrac{(k+1)!}{2\pi i}\int_{\partial D}\dfrac{f(w)}{(w-z)^{(k+1)+1}}dw"

Induction Step

This is our induction step:


"\dfrac{d}{dz}f^{(k)}(z)=\dfrac{k!}{2\pi i}\int_{\partial D}\dfrac{d}{dz}\dfrac{f(w)}{(w-z)^{k+1}}dw="

"=\dfrac{k!}{2\pi i}\int_{\partial D}(k+1)\dfrac{f(w)}{(w-z)^{k+2}}dw="

"=\dfrac{(k+1)!}{2\pi i}\int_{\partial D}\dfrac{f(w)}{(w-z)^{(k+1)+1}}dw=P(k+1)"

So "P(k)=>P(k+1)" and the result follows by the Principle of Mathematical Induction.

There "\forall n\in\N:"


"f^{(n)}(z)=\dfrac{n!}{2\pi i}\int_{\partial D}\dfrac{f(w)}{(w-z)^{n+1}}dw"